OpenStudy (anonymous):

Ok, how can I enter this into a TI-83+ to verify what I did on paper? I am supposed to solve for X without a calc. and then verify it with a calc.... log(3) (x-1) =2

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

log base 3 of x-1?

OpenStudy (anonymous):

I would enter in the y = y =2 and log(x-1)/log(3) then use the intersect for both equations

OpenStudy (anonymous):

or you could enter that same expression and use the table feature to find where it equals 2

OpenStudy (anonymous):

so that would be two separate lines in the Y= screen?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that what you are trying to do? verify using a "graphing" calculator

OpenStudy (zarkon):

are you trying to find the solution with your ti83?

OpenStudy (anonymous):

ok, now if I were to re-write that as a "normal" number, it would be 3^2=(x-1) right?

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

yeah, the instructions say to solve it without a calculator and then verify with a calulator to make sure you did it right

OpenStudy (anonymous):

I came up with X=10 on paper... would you agree?

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

sweet, ok so here is a twist on that question, how would I write 3^2=x-1 in my calculator?

OpenStudy (zarkon):

what is \[\frac{\log(10-1)}{\log(3)}\]

OpenStudy (anonymous):

2

OpenStudy (zarkon):

you just verified your solution

OpenStudy (anonymous):

killer, you are the man!